OpenStudy (anonymous):

integrate 6cos^3(3x)

hartnn (hartnn):

hello :) have you attempted this integral ?

hartnn (hartnn):

there are several ways to solve this integral, one of the ways include, writing \(\cos^3 3x \) as \(\cos^2 x \times \cos x\) and do a 'u' substitution u= sin x

hartnn (hartnn):

will you try this method ? if you get stuck, ask me, i'll help you out :)